When 25.0 mL of 0.25 mol/L LiOH and 25.0 mL of 0.25 mol/L HCl are mixed together, the temperature warms 15.8o C. What is the molar enthalpy of neutralization for LiOH? (3 marks)
1
Expert's answer
2020-09-05T14:42:16-0400

Solution.

"LiOH + HCl = LiCl + H2O"

n(LiOH) = n(HCl)

"m(solution) = V(solution) \times \rho"

m(solution) = 50 g

"Q = c \times m \times \Delta T"

"\Delta H =- Q"

"Q = 4200 \times 0.05 \times 15.8 = 3318 \ J"

"\Delta H = -3318 \ J"

"\Delta H(molar) = \frac{\Delta H}{n(LiOH)}"

"\Delta H(molar) = \frac{-3318}{0.25*0.025} = 530.9 \ kJ/mol"

Answer:

"\Delta H(molar) = 530.9 \ kJ/mol"